OpenStudy (anonymous):

9. Three faces of an ordinary dice are yellow, two faces are red one face is blue. The dice is rolled 3 times. The probability that the yellow red and blue faces appear in the first, second and third throws respectively is

OpenStudy (kropot72):

The outcome of each throw is independent of the outcomes of the other throws. The probabilities are as follows: P(yellow) = 3/6 = 1/2 P(red) = 2/6 = 1/3 P(blue) = 1/6 The probability of getting yellow, red and blue in 3 successive throws is 1/2 * 1/3 * 1/6 There are 3! = 6 permutations of the 3 colors, only one of which has the colors drawn in the required order The probability that the yellow red and blue faces appear in the first, second and third throws respectively is 1/6 * 1/2 * 1/3 * 1/6